OpenStudy (imnotarobot5):

Find the slope of the graph of the relation 2x^3 − 3xy + y^3 = −1 at the point (2, −3).

OpenStudy (photon336):

You could find the derivative of this curve by implicit differentiation. \[\frac{ d }{ dx }(2x^{3}-3(xy)) = \frac{ d }{ dx }(-1)\]

OpenStudy (photon336):

\[6x^{2}-3\frac{ d }{ dx}(xy) = 0 \] \[u = x; v = y, du = 1, dv = y' \] \[6x^{2}-3(y+x \frac{ dy }{ dx }) = 0\]

OpenStudy (photon336):

further manipulation of this and insetting the x and y values will get you the slope. i'll stop here.

OpenStudy (imnotarobot5):

But what about the y^3 @Photon336

OpenStudy (photon336):

I missed that point sorry \[6x^{2}-3(y+x \frac{ dy }{ dx })+3y^{2}(\frac{ dy }{ dx }) = 0\]

OpenStudy (photon336):

what you can do is group terms with dy together

OpenStudy (photon336):

\[6x^{2}-3y+(x+3y^{2})(\frac{ dy }{ dx }) = 0\]

OpenStudy (photon336):

\[3y^{2}-6x^{2} = (x+3y^{2})(\frac{ dy }{ dx })\] now what you can do is solve for dy/dx and plug in your x and y values. originally given.

OpenStudy (photon336):

\[\frac{ 3y^{2}-6x^{2} }{ x+3y^{2} } = \frac{ dy }{ dx }\]

OpenStudy (imnotarobot5):

That makes sense, thank you!